OpenStudy (anonymous):

The bullet travels 148.5 ± 0.5 cm in 5.8 ± 0.1 ms (1 cm = 0.01 m ,1 ms =10−3s). What is the average speed of the bullet in m/s?

ganeshie8 (ganeshie8):

\(\large \text{Average speed} = \frac{\text{distance travelled}}{\text{time taken}} \)

ganeshie8 (ganeshie8):

\(\large \text{Average speed} = \frac{148.5 \pm 0.5 cm}{5.8 \pm 0.1 ms} \)

ganeshie8 (ganeshie8):

\(\large \text{Average speed} = \frac{148.5 }{5.8 } \pm \delta S \)

ganeshie8 (ganeshie8):

\(δS\) is the uncertainty in speed. to get this, you must find the relative uncertainties of distance and time, and add them